OpenStudy (anonymous):

Complete the square, I'm stuck on a step f(x)=-2x^2+12x-8 Then that changes to -2(x^2-6x+9)-8+18 Is it a +18, or -18?

zepdrix (zepdrix):

Plus 18, you need to apply the -2 to your -18 when you pull it out of the brackets.

zepdrix (zepdrix):

\[\Large\rm -2x^2+12x-8\]\[\Large\rm =-2(x^2-6x)-8\]Looks like +9 is what completes the square for us:\[\Large\rm =-2(\color{orangered}{x^2-6x+9}-9)-8\]We subtract 9 at the same time, to keep things balanced. We want to take the -9 out of the brackets, away from our perfect square. So we apply the -2 to it when we take it out.\[\Large\rm =-2(\color{orangered}{x^2-6x+9})-8+18\]I'm assuming that's what you were trying to do, yes? :)

zepdrix (zepdrix):

My first response: `you need to apply the -2 to your -18 when you pull it out of the brackets.` Should have read: `you need to apply the -2 to your -9 when you pull it out of the brackets.`

OpenStudy (anonymous):

On my teacher's answer key at the end she has -8-18, so is that incorrect?

zepdrix (zepdrix):

Lemme double check my work just to make sure c:

zepdrix (zepdrix):

Yah looks like the teacher made a boo boo, it happens :3

zepdrix (zepdrix):

\[\Large\rm -2x^2+12x-8\quad=\quad -2(x-3)^2+10\]

OpenStudy (anonymous):

Thank you